Temperatures of the source and sink between which an ideal heat engine is working are x and y respectively (Kelvin)


Its efficiency is n.


What is its new  efficiency if the source temperature is tripled and sink temperature is doubled?


a)(2n-1)/3


b)(2n+1)/3


c)(3n+1) /2


d)(n+2)/3


 


Plz explain..

Grade:12

1 Answers

bhanu p
12 years ago

We know that  n = 1 - y/x  => n-1 = -y/x ---(1)

similarly, let n' be the new efficiency, and n' = 1 - y'/x' whare y' = 2y and x' = 3x => n' = 1- 2y/3x => n'-1 = -2y/3x ---(2)

(1) / (2) => n-1/n'-1 = 3/2 => n' = (2n + 1)/3 ,therefore option is (b)